# Exploring the Evidence-Based Beliefs of LLM-Based Programming Assistants

### v2

@@ IV-A RQ1: LLM-Generated Beliefs
TABLE II: LLM-Generated Belief (RQ1) Results
Belief Summary ChatGPT GitHub Copilot Gemini Blackbox AI Claude ✓$\sigma^{2}$
@@
C4 Specific experience $>$ general experience N N A Y Y 0.4 0.80
@@
✓[LLM correctness] 0.53 0.53 0.59 0.76 0.76
$\sigma^{2}$ [LLM variance] 0.58 0.68 0.48 0.18 0.18
Finding 1: LLM-based programming assistants agree with most evidence-based claims, but a large number of responses are ambiguous.
IV-B RQ2: LLM-Generated Belief Evidence
We inspected original responses and prompted LLMs to determine the evidence systems used to arrive at their beliefs.
@@ IV-B RQ2: LLM-Generated Belief Evidence
Most responses ($n=64$, 75.2%) lacked clear evidence to support beliefs. We observed 21 instances across three LLMs: Claude ($n=2$), Gemini ($n=2$), and Blackbox AI ($n=17$). For Claude and Gemini, we found generic references—i.e., “research has shown…” (C1, Claude) and “studies on this topic…” (C6, Gemini), without specific references or details. Additionally, one Gemini response included “Research suggests…[1]” (C3), without a citation for “[1]”. Blackbox AI provided the most evidence—totaling 162 sources. Figure 1 presents the types of evidence observed. Most sources were links to SE-related blog posts ($n=46$, 28.4%) and posts on Q&A websites ($n=37$, 22.8%) (i.e., Quora, SE Stack Exchange, and StackOverflow).
@@
001010101020202020303030304040404050505050Tech ReportCode ExampleQ&AResearch PaperTraining ResourceSocial MediaBlog333333333737373731313131262626261616161646464646
Figure 1: Types of Evidence (RQ2) Provided in LLM Belief Responses
@@
We also prompted LLMs to understand evidence for beliefs. Our analysis revealed four high-level categories (see Table III): research ($n=51$), reasoning ($n=49$), training data ($n=48$), and prior response ($n=1$). Most LLMs reported using research to inform their responses. For instance, when describing how it arrived to its response for C1, ChatGPT replied “My response is informed by research and insights gathered from software engineering literature…Studies and articles often discuss how language choice impacts software quality, citing examples and empirical data”. Further, in the prompted response for C3 Gemini incorrectly reports “I included a reference to a relevant study to support the claim that distributed teams can achieve high code quality”. We also observed instances of LLMs using their own reasoning (i.e., “I reached this answer by considering several factors…” (Gemini) and “I recognized the widespread adoption of coding standards in the software industry, which suggests that they are seen as an effective way to improve software quality” (Blackbox AI)) and training data (i.e., “I generate responses based on patterns and information in the data I was trained on” (Copilot)) to support responses. Every prompted response from GitHub Copilot indicated it generates output based on the data it was trained on.
Finding 2: Most LLM-based programming assistants lack evidence for beliefs about SE claims. Only Blackbox AI provided clear evidence, predominantly from blogs and Q&A posts. Responses that reported using research to derive their answers across LLMs were more accurate.
@@ V Implications
LLM-based programming assistants are increasingly used in SE, however our results show these systems lack evidence-based beliefs. Based on these findings, we provide implications for developers adopting LLMs into development workflows.
LLMs are mostly positive, but can be ambiguous on evidence-based practices
We observed the majority of responses from LLMs agreed with SE research claims (see Table II). However, developers utilizing these systems should note they can be noncommittal on important concepts. For instance, four LLM-based programming assistants provided ambiguous responses on the claim well-commented code has fewer defects (C5), providing responses such as “this correlation is not absolute” (ChatGPT). Additional measures should be taken to mitigate ambiguity from LLMs. For example, prior work suggests refining prompts and providing feedback or examples (i.e., to incorporate best practices) can enhance ChatGPT in SE contexts [35].
LLM-generated responses lack credible evidence
While most LLM-based programming assistants claimed to use research, we found a lack of evidence to support their responses. LLMs also frequently used their own reasoning and training data (see Table III), which are unknown to users and can be misleading [51]. Only Blackbox AI consistently provided evidence to support claims made by the generative AI model—but, these were largely from non-scientific sources (see Figure 1). We also observed several complications such as duplicated references and missing citations. The evidence provided by Blackbox AI could contribute to its accuracy in beliefs, as one of the most accurate LLMs. However, the overall lack of evidence should caution practitioners using their output to support development tasks.
LLM beliefs are similar to humans
We conducted a conceptual replication of previous work exploring the beliefs of software engineers at Microsoft [16], substituting LLM-based programming assistants for humans. We do not directly compare our results due to different rating scales (Y/N/A vs. 5-point scale). However, at a glance we found similar trends in beliefs. For instance, the four least controversial human beliefs—on code reviews (C16), coding standards (C15), static analysis tools (C14), and assert statements (C13)—were unanimously agreed on by the LLMs observed in our study. Likewise, three of the most controversial human claims—code quality depends on language (C1), fixing defects is riskier than adding new features (C2), and specific project experience matters more than general experience (C4)—were the most varied among LLMs. This motivates the need for future research to enhance evidence-based SE beliefs and practices for LLMs and human developers.
@@ VI Research Directions
LLMs will continue to innovate SE as adoption increases among developers [5]. As software becomes more complex and widespread and generative AI is integrated into a growing number of SE-focused contexts (i.e., IDEs,666e.g., https://www.jetbrains.com/ai/ Q&A platforms,777e.g., https://stackoverflow.blog/2023/07/27/announcing-overflowai/ etc.), promoting evidence-based SE practices is increasingly critical. Based on our findings, we provide avenues for future research to increase the adoption of evidence-based practices and beliefs among LLMs and the software developers leveraging them to complete SE tasks.
Evidence-Based LLMs
Our results show LLMs lack credible evidence to support claims. To mitigate this, future work can explore training LLMs on evidence-based data. LLM chatbots have been implemented to provide responses backed by scientific evidence to prompts in various domains. For example, Med-PaLM [^8] https://sites.research.google/med-palm/ powers a live Q&A system for medical inquiries about liver cancer [52]. Specialized LLMs have also been developed for technical fields, such as cryptocurrency. [^9] https://app.chaingpt.org/ To promote evidence-based SE, future LLM-based programming assistants could incorporate empirical SE research findings in their training data to provide responses backed by research. In addition, future work can highlight code bases shown to be high-quality and developed using evidence-based tools and practices to incorporate in training processes. This could increase the reliability of LLMs, ideally leading to evidence-based output generated by models.
Evidence-Based LLM Development Workflows
In our study, we observed LLM-based programming assistants can be ambivalent on development practices supported by empirical SE research. Future work can investigate techniques to enhance LLM adoption in SE. For instance, a layered workflow architecture can apply evidence-based practices on LLM-generated code before the output is presented to users. In this context, the layered workflow can incorporate multiple levels of predefined work steps to help organize, control, and execute activities [53]. A potential example could be:
@@
    A user prompts an LLM-based programming assistant to generate code;
@@
    the LLM generates code based on the prompt;
@@
    the generated code and resulting static analysis report are provided to the user to inspect—or, LLM automated program repair capabilities can be leveraged to fix reported bugs before presenting generated code to users.
Prior work uses layers to improve LLMs in SE, such as security within AI/ML pipelines [54]. Recent tools, such as AutoGen, [^10] https://www.microsoft.com/en-us/research/project/autogen/ also provide multi-agent support for customized workflows. Future work can explore infrastructure to augment LLM-based development workflows by incorporating research-based tools and practices. Studies can also investigate the effects of comparable approaches on code quality and developer behavior.
LLMs for Translating Research Evidence
Empirical SE research aims to provide evidence supporting effective software development [7]. However, translating research findings into practice is challenging [9]. Horvitz et al. suggest one of the top priorities for the future of AI is to power resources to bridge the academia-industry gap [55]. Thus, future work can explore using LLMs to summarize and distribute SE research findings to practitioners. While our findings show zero-shot prompting of LLMs regarding evidence-based practices often provides ambiguous and unsupported responses, prior work suggests LLMs are effective for summarizing natural language text and code [56]. Previous attempts to convey human summaries of research findings have shown to be ineffective [57]. However, future work can explore the potential of LLM-generated summaries and the role they can play in supporting the behavior of software practitioners.
@@ VII Conclusion
To support software development, empirical research has offered a variety of tools and practices to enhance the development and maintenance of software. More recently, LLMs have transformed the development and maintenance of software—receiving rapid adoption for their ability to automatically support a wide variety of SE-related tasks. However, little is known about these AI-driven systems with regard to evidence-based practices purported by empirical SE research. We analyze the beliefs of popular LLM-based programming assistants through a conceptual replication of prior work [16], and explore the evidence for these beliefs. Our findings show LLMs do not provide a clear stance on many practices and lack credible evidence. Based on these emerging results, we provide implications and future research directions to enhance empirically-based decision making with LLMs in SE contexts.

### v4

@@ IV-A RQ1: Generative AI-Generated Beliefs
TABLE II: Generative AI-Generated Belief (RQ1) Results
Belief Summary ChatGPT GitHub Copilot Gemini Blackbox AI Claude ✓% $\sigma^{2}$
@@
C4 Specific experience >> general experience N N A Y Y 0.4 0.80
@@
✓% [Generative AI tool correctness] 0.53 0.53 0.59 0.76 0.76
$\sigma^{2}$ [Generative AI tool variance] 0.58 0.68 0.48 0.18 0.18
Finding 1: Generative AI tools agree with most evidence-based claims, but many responses are ambiguous.
IV-B RQ2: Generative AI-Generated Belief Evidence
We inspected original responses and prompted AI tools to determine the evidence systems used to arrive at their beliefs.
@@ IV-B RQ2: Generative AI-Generated Belief Evidence
Most responses ($n=64$, 75.2%) lacked clear evidence to support beliefs. We observed 21 instances across three generative AI tools: Claude ($n=2$), Gemini ($n=2$), and Blackbox AI ($n=17$). For Claude and Gemini, we found generic references—i.e., “research has shown…” (C1, Claude) and “studies on this topic…” (C6, Gemini), without specific references or details. Additionally, one Gemini response included “Research suggests…[1]” (C3), without a citation for “[1]”. Blackbox AI provided the most evidence—totaling 162 sources. Figure 1 presents the types of evidence observed. Most sources were links to SE-related blog posts ($n=46$, 28.4%) and posts on Q&A websites ($n=37$, 22.8%) (i.e., Quora, SE Stack Exchange, and StackOverflow).
@@
010102020303040405050Tech ReportCode ExampleQ&AResearch PaperTraining ResourceSocial MediaBlog333337373131262616164646
Figure 1: Types of Evidence (RQ2) Provided in Generative AI Belief Responses
@@
We also prompted generative AI tools to understand evidence for beliefs. Our analysis revealed four high-level categories (see Table III): research ($n=51$), reasoning ($n=49$), training data ($n=48$), and prior response ($n=1$). Most tools reported using research to inform their responses. For instance, when describing how it arrived to its response for C1, ChatGPT replied “My response is informed by research and insights gathered from software engineering literature…Studies and articles often discuss how language choice impacts software quality, citing examples and empirical data”. Further, in the prompted response for C3 Gemini incorrectly reports “I included a reference to a relevant study to support the claim that distributed teams can achieve high code quality”. We also observed instances of AI tools using their own reasoning (i.e., “I reached this answer by considering several factors…” (Gemini) and “I recognized the widespread adoption of coding standards in the software industry, which suggests that they are seen as an effective way to improve software quality” (Blackbox AI)) and training data (i.e., “I generate responses based on patterns and information in the data I was trained on” (Copilot)) to support responses. Every prompted response from GitHub Copilot indicated it generates output based on the data it was trained on.
Finding 2: Most generative AI tools lack evidence for SE claims. Only Blackbox AI provided clear evidence, predominantly from blogs and Q&A posts.
@@ V Implications
Generative AI tools are increasingly used to support SE tasks. However, our results show these systems lack evidence-based beliefs—potentially inhibiting their utility for practical software development. Based on these findings, we provide implications for developers integrating generative AI tools and LLMs into development workflows.
Generative AI is mostly positive, but can be ambiguous on evidence-based practices
We observed the majority of responses from generative AI tools agreed with SE research claims (see Table II). However, developers utilizing these systems should note they can be noncommittal on important concepts. For instance, four generative AI tools provided ambiguous responses on the claim well-commented code has fewer defects (C5), providing responses such as “this correlation is not absolute” (ChatGPT). Additional measures should be taken to mitigate ambiguity from AI tools. For example, prior work suggests refining prompts and providing feedback or examples (i.e., to incorporate best practices) can enhance ChatGPT in SE contexts [37].
AI-generated responses lack credible evidence
While most generative AI tools claimed to use research, we found a lack of evidence to support their responses. AI tools also frequently used their own reasoning and training data (see Table III), which are unknown to users and can be misleading [53]. Only Blackbox AI consistently provided evidence to support claims made by the generative AI model—but, these were largely from non-scientific sources (see Figure 1). We also observed several complications such as duplicated references and missing citations. The evidence provided by Blackbox AI could contribute to its accuracy in beliefs, as one of the most accurate AI tools. However, the overall lack of evidence forming generative AI tool beliefs should caution practitioners using their output to support development tasks.
Generative AI beliefs are similar to humans
We conducted a conceptual replication of previous work exploring the beliefs of software engineers at Microsoft [17], substituting generative AI tools for humans. We do not directly compare our results due to different rating scales (Y/N/A vs. 5-point scale). However, at a glance we found similar trends in beliefs. For instance, the four least controversial human beliefs—on code reviews (C16), coding standards (C15), static analysis tools (C14), and assert statements (C13)—were unanimously agreed on by the AI tools observed in our study. Likewise, three of the most controversial human claims—code quality depends on language (C1), fixing defects is riskier than adding new features (C2), and specific project experience matters more than general experience (C4)—were the most varied among AI tools. This motivates the need for future research to enhance evidence-based beliefs and practices for generative AI-based SE tools and human developers.
@@ VI Research Directions
Generative AI tools will continue to innovate software development. As software becomes more complex and widespread, AI adoption is increasing among developers [6]. Generative AI is also increasingly integrated into a growing number of SE-focused contexts (i.e., IDEs,666e.g., https://www.jetbrains.com/ai/ Q&A platforms,777e.g., https://stackoverflow.blog/2023/07/27/announcing-overflowai/ etc.). Thus, promoting evidence-based SE practices is increasingly critical. Based on our findings, we provide avenues for future research to increase the adoption of evidence-based practices and beliefs among generative AI tools and the software developers leveraging them to complete SE tasks.
Evidence-Based Generative AI
Our results show AI lacks credible evidence to support claims. To mitigate this, future work can explore training generative AI on evidence-based data. AI-based chatbots have been implemented to provide responses backed by scientific evidence to prompts in various domains. For example, Med-PaLM [^8] https://sites.research.google/med-palm/ powers a live Q&A system for medical inquiries about liver cancer [54]. Specialized models have also been developed for technical fields, such as cryptocurrency. [^9] https://app.chaingpt.org/ To promote evidence-based SE, future generative AI tools could incorporate empirical SE research findings in their training data to provide responses backed by research. In addition, future work can highlight code bases shown to be high-quality and developed using evidence-based tools and practices to incorporate in training processes. This could increase the reliability of AI, ideally leading to evidence-based output generated by models.
Evidence-Based AI Development Workflows
In our study, we observed generative AI tools can be ambivalent on development practices supported by empirical SE research. Future work can investigate techniques to enhance AI adoption in SE. For instance, a layered workflow architecture can apply evidence-based practices on AI-generated code before the output is presented to users. In this context, the layered workflow can incorporate multiple levels of predefined work steps to help organize, control, and execute activities [55]. A potential example could be:
@@
    A user prompts a generative AI tool to generate code;
@@
    the AI tool generates code based on the prompt;
@@
    the generated code and resulting static analysis report are provided to the user to inspect—or, AI-based automated program repair capabilities can be leveraged to fix reported bugs before presenting generated code to users.
Prior work uses layers to improve generative AI integration in SE, such as security within AI/ML pipelines [56]. Recent tools, such as AutoGen, [^10] https://www.microsoft.com/en-us/research/project/autogen/ also provide multi-agent support for customized workflows. Future work can explore infrastructure to augment generative AI-based development workflows by incorporating research-based tools and practices. Studies can also investigate the effects of comparable approaches on code quality and developer behavior.
Generative AI for Translating Research Evidence
Empirical SE research aims to provide evidence supporting effective software development [8]. However, translating research findings into practice is challenging [10]. Horvitz et al. suggest one of the top priorities for the future of AI is to power resources to bridge the academia-industry gap [57]. Thus, future work can explore using generative AI to summarize and distribute SE research findings to practitioners. While our findings show zero-shot prompting regarding evidence-based practices often provides ambiguous and unsupported responses, prior work suggests LLMs are effective for summarizing natural language text and code [58]. Previous attempts to convey human summaries of research findings have shown to be ineffective [59]. However, future work can explore the potential of AI-generated summaries and the role they can play in supporting the behavior of software practitioners.
@@ VII Conclusion
To support software development, empirical research offers a variety of tools and practices to enhance software development. Recently, generative AI tools have transformed the development and maintenance of software—receiving rapid adoption for their ability to automatically support a wide variety of SE-related tasks. However, little is known about these LLM-driven systems with regard to evidence-based practices purported by empirical SE research. We analyze the beliefs of popular generative AI tools used for development tasks through a conceptual replication of prior work [17], and explore the evidence for these beliefs. Our findings show generative AI tools often do not provide a clear stance on SE practices, such as the benefits of well-commented code, and lack credible evidence. Based on these results, we provide implications and future research directions to enhance evidence-based decision making with generative AI in SE contexts.